Priscilla Diller quit her $60,000 per year job with South Central Construction to open her
own company, Diller Didactic Consulting. She rented an office in the Peach Street
Professional Building for $550 a month, hired a receptionist for $1,500 a month, and
leased a computer from Silly Clone Systems for $130 per month.
At first, business was slow, so Priscilla paid $800 to use a mailing list and sent out 1,200
advertising brochures. The cost of postage and printing was $1.37 per brochure. Business
improved immediately, and no further advertising has been necessary.
Priscilla's business improved so much that she hired two associates for $5,000 per month
and installed them in the remaining two offices in the Peach Street Professional Building,
and then she bought the building for $198,000. She replaced her one leased computer
with three computers of the same type that she purchased from Silly Clone Systems for
$22,000 and she talked
1
Expert's answer
2021-08-03T13:03:22-0400

1) implicit cost=60000

Explicit cost=rent+ receptionist cost+computer cost+mailing cost+ advertising cost

"550+1500+130+800+(1.37)×290=\$3254"

Total economic cost=explicit cost+implicit cost"=3254+60000=\$63254"


2)when business improved

Explicit cost "3254+2×(5000)+198000+22000=224254+60000=\$284254"